OpenStudy (anonymous):

4-5(a+2)=3(a+1)=1 someone help?

OpenStudy (saifoo.khan):

Turing will help! :D

OpenStudy (anonymous):

-.-

OpenStudy (saifoo.khan):

Chill.

OpenStudy (anonymous):

i am chill.

OpenStudy (saifoo.khan):

really?

OpenStudy (anonymous):

YES REALLY!

OpenStudy (turingtest):

you seem to have an extra = sign, perhaps it should be 4-5(a+2)=3(a+1)-1 or something?

OpenStudy (anonymous):

on my paper it has 2 equal signs o:

OpenStudy (turingtest):

This leads to a direct contradiction, so there is a typo

OpenStudy (anonymous):

The a equal -1

OpenStudy (anonymous):

o: so what if its - instead? lol

OpenStudy (anonymous):

a's

OpenStudy (anonymous):

do the distributive property so: 4-5a+10=3a+3=1 can u do the rest?

OpenStudy (anonymous):

i had this same promblem there is only one equal sign

OpenStudy (turingtest):

^that makes sense

OpenStudy (turingtest):

@christinaxxx I can't, can you?

OpenStudy (turingtest):

I'm going to assume it is 4-5(a+2)=3(a+1)-1 4-5a-10=3a+3-1 8a=-8 a=-1 which would give applepie's answer, so that is probably what the problem is supposed to be

OpenStudy (anonymous):

im confused!

OpenStudy (anonymous):

so it would be -1?

OpenStudy (turingtest):

IF the equation is as I typed it, then yes. If it is as you typed it there is no solution.